Given the following LP problem formulation and output data, perform the analysis below.

Max.    100X1 + 120X2 + 150X3 + 125X4

s.t        X1 + 2X2 + 2X3 + 2X4 < 108            (C1)

           

            3X1 + 5X2 + X4 < 120                       (C2)

            X1 + X3 < 25                                      (C3)

X2 + X3 + X4 > 50                             (C4)

OPTIMAL SOLUTION:

Objective Function Value = 7475.000

Variable

Value

Reduced Costs

X1

8.000

0.500

X2

0.000

5.000

X3

17.000

0.000

X4

*A

0.000

Constraint

Slack / Surplus

Dual Prices

1

0.000

*B

2

63.000

0.000

3

*C

25.000

4

0.000

-25.000

            Objective Co-efficient Ranges

Variable

Lower Limit

Current Value

Upper Limit

X1

87.500

100.000

No Upper Limit

X2

No lower limit

120.000

125.000

X3

125.000

150.000

162.500

X4

120.000

125.000

150.000

            Right Hand Side Ranges

Constraint

Lower Limit

Current Value

Upper Limit

1

100.000

108.000

123.750

2

57.000

120.000

No Upper Limit

3

8.000

25.000

58.000

4

41.000

50.000

54.000

  1. Find *A
  1. How much of constraint C2 is used?
  1. In each of the following, calculate the resulting change in the objective function, if possible and show you calculations. If not possible, explain why not.

  1. The obj.func.co-eff of X4 increases from 125 to 135.
  1. The level of C4 decreases from 50 to 42.
Answer #1

Question – a:

The objective function is: 100X1 + 120*X2 + 150*X3 + 125*X4

Here,

X1 = 8

X2 = 0

X3 = 17

X4 = *A

Objective function value = 7475

Putting values in equation:

100*8 + 120*0 + 150*17 + 125*(*A) = 7475

800 + 2550 + 125*(*A) = 7475

125*(*A) + 3350 = 7475

125*(*A) = 7475 – 3350 = 4125

(*A) = 4125/125 = 33

Answer is: *A = 33

Question – b:

Constraint 2 is: 3X1 + 5x2 + X4

Putting values in equation:

3*8 + 0 + 33 = 24 + 33 = 57

Hence, 57 units of Constraint 2 have been used.

Question – c:

1. As the objective coefficient upper limit is 150, increasing it from 125 to 135 wont affect the optimal solution. The current solution will remain optimal.

Objective function will be: 100X1 + 120X2 + 150X3 + 135X4

X1 = 8, X2 = 0, X3 = 17, X4 = 33

Objective function value = 100*8 + 120*0 + 150*17 + 135*33 = 7805

2. The current solution will remain optimal by decreasing the C4 from 50 to 42 as the lower limit is 41. But we wont be able to calculate the objective function value as the shadow price has not been provided.

********

IF YOU HAVE ANY DOUBT, KINDLY COMMENT. I WOULD LOVE TO HELP YOU!!!

IF YOU LIKE THE ANSWER, PLEASE GIVE AN UP-VOTE OR THUMB UP. THIS WILL ENCOURAGE ME TO ANSWER MORE!!